# SSVQ: Unleashing the Potential of Vector Quantization with Sign-Splitting

### v1

@@ 5.1 Experimental setup
Implementation details. We employ the AdamW optimizer with a cosine annealing learning rate schedule across all tasks. To ensure reproducible comparative analysis, we maintain a consistent 10-epoch training regimen without distillation for both classification and detection tasks, while reproducing uniform quantization benchmarks under the standard w2 configuration. Image generation experiments follow the established pipeline in  [15], while semantic segmentation builds upon  [1]. The sign variation ratio is modulated through either the learning rate or $\theta$. Detailed hyperparameter specifications are available in the supplementary material.
Comparsion methods. Given the limited availability of VQ implementations validated on diverse modern vision architectures (many previous VQ works [34, 24, 17, 5, 16, 44] still focus greatly on ResNets), we systematically implemented conventional VQ with identical training protocols to SSVQ for each task category, establishing comprehensive baseline evaluations. For comparative analysis, we present discrete data points from contemporary VQ methods (MVQ [16], HQ [17], DKM [5]) and state-of-the-art UQ approaches (LSQ [6], PvQ [13], Osqat [27], OFQ [18], NME [15], PCR [36], Quest [41], MPQ-DM [7]).
@@ 5.2 Performance Evaluation
Models MobileNet-v2 EfficientNet-lite
Methods MVQ[16] HQ[17] SSVQ MVQ[16] SSVQ
@@
Method VQ PvQ [14] MVQ [16] SSVQ
C.R 17.5$\times$ 16 $\times$ 19$\times$ 19$\times$
mIoU 65.7 19.2 66.5 67.7
@@
Table 5: Quantization results of object detection and instance segmentation on COCO dataset. ’bbox AP’ is the bounding box average precision for object detection, and ’mask AP’ is the mask average precision for instance segmentation.
@@
COCO Prompts FP: 31.4/ 31.3* PCR [36] 2 311.0$\dagger$ 19.1 $\dagger$
NME [15] 2 173.4$\dagger$ 18.3$\dagger$
@@
SD Prompts FP:31.6 PCR [36] 2 322.6$\dagger$ 19.4$\dagger$
NME [15] 2 102.8$\dagger$ 22.3$\dagger$
@@
COCO Prompts FP:31.5 PCR [36] 2 330.8$\dagger$ 19.2$\dagger$
NME [15] 2 137.6$\dagger$ 23.6$\dagger$
@@
SD Prompts FP:31.6 PCR [36] 2 364.6$\dagger$ 20.1$\dagger$
NME [15] 2 181.7$\dagger$ 21.3$\dagger$
@@
COCO Prompts FP:31.7 NME [15] 2 214.9$\dagger$ 21.5$\dagger$
VQ 2 16.8 30.6
SSVQ 2 12.8 31.7
SD Prompts FP:30.6 NME [15] 2 266.9$\dagger$ 20.8$\dagger$
VQ 2 23.4 29.7
SSVQ 2 17.4 30.7
Table 6: Quantization results of image generation of Stable Diffusion models on COCO 2017 val prompts and SD prompts. We generate 5000 images for each prompt set. CLIP scores were computed using ViT-B/16, with FP model baselines shown below prompts. $\dagger$ means reproduced results. $\downarrow$ means lower is better. $\dagger$ means reproduced results. * means validation on COCO2014 val prompts.
@@ 6 System Evaluation
Figure 7: Arthitecture overview of our hardware simulator supporting SSVQ.
@@ 6.2 Architecture support for SSVQ
To optimize the utilization of main memory bandwidth during lookup table operations, it is essential to ensure that the decoding speed aligns with the bandwidth of the main memory. When reconstructing a weight subvector of size $d$, $\log_{2}{k}$ index bit and $d$ sign bit is required. Denoting the main memory bandwidth as $B$ (bit/cycle), we must guarantee that the decoding parallelism $N$ satisfies
$N>\frac{B}{Ceil_{8}(\log_{2}{k}+d)}$ (9)
where $Ceil_{8}$ is the function that rounds a number up to the nearest multiple of 8. To accomplish this, We implement the codebook using a multi-read-write-port register file (Regfile). In our case, we set 2 codebooks each with 4 read ports.
@@ 9.1 Comparison on visual similarity
COCO Prompts NME [15] 0.66 0.39 11.5
VQ 0.58 0.45 12.7
SSVQ 0.51 0.49 13.3
SD Prompts NME [15] 0.72 0.30 10.9
VQ 0.63 0.46 13.2
SSVQ 0.53 0.48 13.9
@@
COCO Prompts NME [15] 0.72 0.38 9.6
VQ 0.61 0.43 10.3
SSVQ 0.58 0.46 10.7
SD Prompts NME [15] 0.75 0.28 6.80
VQ 0.64 0.33 9.10
SSVQ 0.60 0.37 9.57
@@
To provide a comprehensive evaluation of generation consistency, we introduce visual similarity metrics as complementary assessment criteria. Specifically, we employ three widely-adopted measures: SSIM (structural similarity), LPIPS (learned perceptual image patch similarity), and PSNR (peak signal-to-noise ratio), maintaining consistency with the evaluation protocol used in NME [15]. As demonstrated in Tab. 9, our SSVQ framework achieves an average improvement of 10% in visual similarity metrics compared to the standard VQ baseline, highlighting its enhanced capability in preserving visual fidelity.
@@ 9.2 Visual results
9.3 Hyper parameters
Models batch size lr lr_mask epochs $\theta$ Optimizer
@@ 9.3 Hyper parameters
DeiT-Tiny 256 2.5e-4 2.5e-4 10 4 AdamW

### v2

@@ 5.2 Experimental setup
Implementation details. We employ the AdamW optimizer with a cosine annealing learning rate schedule across all tasks. To ensure reproducible comparative analysis, we maintain a consistent 10-epoch training regimen without distillation for both classification and detection tasks. Uniform quantization benchmarks are reproduced under the standard ‘w2‘ configuration. Image generation experiments follow the established pipeline in  [18], while semantic segmentation builds upon  [1]. The sign variation ratio is modulated through either the learning rate or the initialization parameter $\alpha$. Detailed hyperparameter specifications are available in the supplementary material.
Comparison methods. Given the limited availability of VQ implementations validated on diverse modern vision architectures (many previous VQ works [39, 28, 20, 6, 19, 50] still focus greatly on ResNets), we systematically implemented conventional VQ baseline with identical training protocols to SSVQ for each task category, establishing comprehensive baseline evaluations. For comparative analysis, we also present results from contemporary VQ methods (MVQ [19], HQ [20], DKM [6]) and state-of-the-art UQ approaches (LSQ [7], PvQ [16], Osqat [31], OFQ [21], EMF [18], PCR [41], Quest [47], MPQ-DM [8]).
@@ 5.3 Performance Evaluation
Models MobileNet-v2 (71.9) EfficientNet-lite (75.4)
Methods MVQ HQ SSVQ MVQ SSVQ
@@
Method FP VQ PvQ [17] MVQ [19] SSVQ
C.R - 17.5$\times$ 16 $\times$ 19$\times$ 19$\times$
mIoU 72.9 65.7 19.2 66.5 67.7
@@
Table 5: Quantization results of object detection and instance segmentation on COCO dataset. ’bbox AP’ is the bounding box average precision for object detection, and ’mask AP’ is the mask average precision for instance segmentation. $\dagger$ denotes our reproduced results.
@@
COCO
Prompts
FP 16 0 31.4
PCR [41] 2 311.0$\dagger$ 19.1 $\dagger$
EMF [18] 2 173.4$\dagger$ 18.3$\dagger$
@@
SD
Prompts
FP 16 0 31.6
PCR [41] 2 322.6$\dagger$ 19.4$\dagger$
EMF [18] 2 102.8$\dagger$ 22.3$\dagger$
@@
COCO
Prompts
FP 16 0 31.5
PCR [41] 2 330.8$\dagger$ 19.2$\dagger$
EMF [18] 2 137.6$\dagger$ 23.6$\dagger$
@@
SD
Prompts
FP 16 0 31.6
PCR [41] 2 364.6$\dagger$ 20.1$\dagger$
EMF [18] 2 181.7$\dagger$ 21.3$\dagger$
@@
COCO
Prompts
FP 16 0 31.7
EMF [18] 2 214.9$\dagger$ 21.5$\dagger$
VQ 2 18.5 30.6
SSVQ 2 11.2 31.7
SD
Prompts
FP 16 0 30.6
EMF [18] 2 266.9$\dagger$ 20.8$\dagger$
VQ 2 28.2 29.7
SSVQ 2 14.2 30.7
Table 6: Quantization results of image generation of Stable Diffusion models on COCO 2017 val prompts and SD prompts. We generate 5000 images for each prompt set. CLIP scores were computed using ViT-B/16, with FP model baselines shown below prompts. $\dagger$ means reproduced results. $\downarrow$ means lower is better. * means validation on COCO2014 val prompts.
@@
Results for NLP tasks. To further assess the generalization of our method, we validate it on Llama3.2-1B, using hessian-weighted k-means[15] as initialization. Results in Tab.7 clearly show SSVQ’s superiority over VQ.
[…]
FP 16 9.7 12.1 0.60 0.65 0.31
VQ 2.5 20.8 24.9 0.53 0.50 0.23
SSVQ 2.5 11.9 15.4 0.57 0.61 0.29
VQ 2 63.1 58.1 0.51 0.40 0.21
SSVQ 2 13.9 17.5 0.54 0.55 0.27
Table 7: Comparison between VQ and SSVQ on Llama3.2-1B. We report PPL on Wiki2 & C4, and accuracy for zero-shot tasks.
@@ 6 System Evaluation
Figure 7: Architecture overview of our hardware simulator supporting SSVQ.
@@ 7 Conclusion
8 Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant No. 62274142), Sino-German Mobility Programme (Grant No. M-0499), and Zhejiang University - Vivo Information Technology Joint Research Center.
@@ 8 Acknowledgements
9 Analysis of initial clustering error
Figure 9: Comparison of initial clustering error between VQ and SSVQ.
Although storing the sign bit requires an additional 1 bit, our method does not introduce extra storage overhead. After extracting the sign bits, we can cluster the absolute values of the weights. Naturally, the similarity among non-negative weights is higher, so we only need fewer bits for cluster indexing. In Fig. 9, we compare the initial clustering errors of SSVQ and VQ. The results show that under the same compression rate, their clustering errors are comparable, which validates our hypothesis.
@@ 10 Further analysis with dimension
11 Effect of Hessian weighted k-means
SqueezeLLM [15] employs Hessian-weighted initialization. Specifically, its objective function is designed to minimize the final task loss caused by quantized weights
$\delta L=L(W)-L(\hat{W})$ (10)
We can achieve this objective through Hessian matrix approximation. Assuming that $\hat{W}$ represents a small perturbation on W, we perform a second-order Taylor expansion on Eq. 10 to obtain:
$\delta L\approx g^{T}(W-\hat{W})+\frac{1}{2}(W-\hat{W})^{T}H(W-\hat{W})$ (11)
Here, $g$ is the gradient, and $H$ is the Hessian matrix of the weights (the second derivative of the Loss with respect to the weights). Assuming the pretrained model has converged, the gradient can be approximated as zero. Furthermore, we can approximate $H$ by retaining only its diagonal elements (which typically dominate).
$\delta L\approx\sum_{i}H_{ii}\|\mathbf{w}_{i}-\mathbf{c}_{A(i)}\|^{2}$ (12)
We observed that this approach significantly benefits LLMs, but fails to demonstrate clear advantages over standard k-means on vision-related tasks, especially after fine-tuning. Thus, we retain standard k-means in other experiments.
12 Detailed quantization settings.
In this section, we listed the detailed quantization settings of our experiments, specifically $k$ and $d$. Classification tasks (Tab. 10), detection and segmentation tasks (Tab. 11), image generation and nlp tasks (Tab. 12).
SSVQ, d=8
k 8 16 32 64 128
MobileNet-v2 22.9 20.7 18.6 16.6 -
MobileNet-v3 23.2 21.3 19.6 18.1 16.7
EfficientNet-lite0 23.0 20.9 19.0 17.2 -
ConvNext-Tiny 23.2 21.3 19.6 18.1 16.8
DeiT-Tiny 23.1 21.1 19.2 17.5 15.8
DeiT-Small 23.2 21.3 19.6 18.1 -
VQ, d=4
k 32 64 128 256 -
MobileNet-v2 - 20.1 16.6 - -
MobileNet-v3 23.5 18.7 - - -
EfficientNet-lite0 - 20.6 17.2 - -
ConvNext-Tiny - 21.2 18.1 - -
DeiT-Tiny - 20.8 17.5 - -
DeiT-Small - 21.2 18.1 - -
VQ, d=8
k 128 256 512 1024 2048
MobileNet-v2 - - 20.9 16.2 -
MobileNet-v3 19.8 - - - -
EfficientNet-lite0 - 17.8 - -
ConvNext-Tiny - - 25.6 21.6 18.1
DeiT-Tiny - 24.7 18.6 - -
DeiT-Small - - - 25.1 20.7
Table 10: $k$&$d$ settings and corresponding compression ratios of classification tasks (Fig. 5). Notably, results with compression ratios $\leq$16× or unacceptably low accuracy are omitted from the figure.
Models Methods k d C.R.
Yolo-v9 VQ 64 4 19.67
SSVQ 16 8 20.47
Gelan-C VQ 64 4 20.98
SSVQ 16 8 21.15
Table 11: $k$&$d$ settings and corresponding bits/weight of detection and segmentation tasks (Tab. 5).
Methods k d Index bit Mask bit Total bit
VQ 1024 4 2.5 0 2.5
VQ 256 4 2 0 2
SSVQ 64 4 1.5 1 2.5
SSVQ 256 8 1 1 2
Table 12: $k$ and $d$ settings and corresponding bits/weight of generation (Tab. 6) and NLP tasks (Tab. 7).
@@ 13.1 Comparison on visual similarity
COCO Prompts EMF [18] 0.66 0.39 11.5
VQ 0.59 0.44 12.7
SSVQ 0.48 0.50 13.8
SD Prompts EMF [18] 0.72 0.30 10.9
VQ 0.64 0.46 13.3
SSVQ 0.52 0.48 14.4
@@
COCO Prompts EMF [18] 0.72 0.38 9.6
VQ 0.59 0.42 10.4
SSVQ 0.54 0.48 11.0
SD Prompts EMF [18] 0.75 0.28 6.80
VQ 0.64 0.34 9.40
SSVQ 0.58 0.40 9.90
@@
To provide a comprehensive evaluation of generation consistency, we introduce visual similarity metrics as complementary assessment criteria. Specifically, we employ three widely-adopted measures: SSIM (structural similarity), LPIPS (learned perceptual image patch similarity), and PSNR (peak signal-to-noise ratio), maintaining consistency with the evaluation protocol used in EMF [18]. As demonstrated in Tab. 13, our SSVQ framework achieves an average improvement of 10% in visual similarity metrics compared to the standard VQ baseline, highlighting its enhanced capability in preserving visual fidelity.
@@ 13.2 Visual results
Models batch size lr lr_mask epochs $\alpha$ Optimizer
@@
DeiT-Small 256 2.5e-4 2.5e-4 10 4 AdamW